OpenStudy (anonymous):

What is the equation of the circle that has a center of (-3.-6) and passes through (-4,8)?

OpenStudy (anonymous):

firstly, do you know the distance formula?

OpenStudy (anonymous):

Yepppp i know thr equation of a circle as well but i dont know how to solve that problem

OpenStudy (anonymous):

so whats your distance from these two points?

OpenStudy (anonymous):

Hold on

OpenStudy (anonymous):

Square root of 196

OpenStudy (anonymous):

you sure its not 197?

OpenStudy (anonymous):

sqrt root of 197

OpenStudy (anonymous):

Yes thts wa i meant sorry

OpenStudy (anonymous):

sweet

OpenStudy (anonymous):

so do you know the general notation for an equation of a circle?

OpenStudy (anonymous):

Equation of a Circle The equation of the circle with center located at (a,b) and with radius r is r2=(x−a)2+(y−b)2 Note that the radius squared is on the left-hand side of the equation.

OpenStudy (anonymous):

r^2=(x-a)^2+(y-b)^2

OpenStudy (anonymous):

Yep

OpenStudy (anonymous):

so what do you think the equation of the circle is with this information

OpenStudy (helder_edwin):

the radius is the distance from the center of the circle to ANY of the points over the circle

OpenStudy (helder_edwin):

this means that if C(-3,-6) and P(-4,8) is a point beloging to the circle then \[\large r=d(C,P) \]